OpenStudy (anonymous):

let f(x)=x^3+ax^2+2a^2x+a such that f(x) has a point of inflection located at x=2. What is the value of a?

OpenStudy (rational):

Hint : a necessary condition for inflection point is, the "second derivative must be 0"

OpenStudy (anonymous):

Ok so would the derivative be f'(x)=3x^2+6ax+a?

OpenStudy (rational):

f(x)=x^3+ax^2+2a^2x+a f'(x) = 3x^2 + 2ax + 2a^2 f''(x) = 6x + 2a

OpenStudy (rational):

set f''(2) = 0 and solve "a"

OpenStudy (anonymous):

ok so a is -6?

OpenStudy (rational):

Yep!

OpenStudy (anonymous):

nice! thanks a lot!

OpenStudy (rational):

yw